# Small-Signal Model for Inhomogeneous Helix Traveling-Wave Tubes using Transfer Matrices

### v1

@@ V Interaction Impedance and Attenuation Coefficient of the SWS
where $\beta_{\mathrm{c}}$ is the phase constant within the fundamental Brillouin zone, defined here as $\beta_{\mathrm{c}}d/\pi\in\left[-1,1\right]$. For a helix-type TWT, the interaction impedance is typically evaluated within the fundamental Brillouin zone, where the beam-wave interaction is desired to occur. However, for space harmonic SWSs, such as the serpentine waveguide or coupled-cavity structure, one must consider higher order space harmonics when determining interaction impedance. Furthermore, $\left|E_{z}\left(\beta_{\mathrm{c}}\right)\right|$ is the magnitude of the axial electric field phasor (i.e., the $z$-component), for a given phase constant, along the center of the SWS where the electron beam will be introduced, and $P\left(\beta_{\mathrm{c}}\right)$ is the time-average total power flux of the mode propagating through the SWS (Gewartowski and Watson, 1965, (Ch. 10)). The quantity $\left|E_{\mathrm{z}}\left(\beta_{\mathrm{c}}\right)\right|$ is calculated from full-wave eigenmode simulations as
@@
as has also been explained in (MM, Aditya, and Chua, 2017; de Alleluia et al., 2020; Marosi et al., 2022). Conveniently, the EM energy simulated within the unit cell between periodic boundaries in the eigenmode solver of CST Studio Suite is always normalized to 1 Joule for each eigenmode solution. Thus, the time average power flux may be simply calculated as $P=\left(1\leavevmode\nobreak\ \mathrm{Joule}\right)v_{\mathrm{g}}/d$, where the group velocity $v_{\mathrm{g}}=d\omega/d\beta_{\mathrm{c}}$ can be determined directly from the dispersion diagram using numerical differentiation. Note that the eigenmode solver of CST Studio Suite neglects losses when finding the wavenumbers and EM fields of eigenmodes, so the electric field and wavenumbers used for computing the interaction impedance also neglect the effect of losses.
@@ VI.1 Single-Stage TWT Transfer Matrix
Our first example structure, the single stage TWT, is composed of a uniform-pitch helix SWS with $N_{\mathrm{c}}=95$ turns (SWS length $l=N_{\mathrm{c}}d=98.8$ mm), with unit cell dimensions shown in Fig. 1 and described in Section II. Note again that, in our examples, the number of unit cells, $N_{\mathrm{c}}$, is different from the number of segments, $S$, that the TWT stage is divided into to make the model more flexible. Based on the specific loss profiles shown in in Appendix C, we implement a nonuniform position-dependent Gaussian-shaped loss pattern in our full-wave PIC simulations by setting a loss tangent for the dielectric support rods in each individual segment (with segment length $\Delta l$, as described above) along the length of the SWS, as explained in Appendix C. At the port-ends of the SWS, the loss pattern for the single-stage SWS has a frequency-dependent attenuation coefficient of $\alpha_{\mathrm{c}}=\alpha_{\mathrm{min}}$ (calculated using Eqn. (7)) . At the midpoint between the input and output ports (i.e., at $z=l/2$) the SWS has a frequency-dependent peak attenuation coefficient $\alpha_{\mathrm{max}}=80\alpha_{\mathrm{min}}$ (determined by scaling the dielectric loss tangent by a dimensionless factor of $\tau=2000$) , and an effective loss pattern length of $l_{\alpha}=30$ mm, equal to the FWHM of the Gaussian function, as is explained in Appendix C. The quantities $\alpha_{\mathrm{min}}$ and $\alpha_{\mathrm{max}}$ correspond the attenuation coefficients in the “clean” and “deep” regions of the loss-coated dielectric support rods, respectively (i.e., the “clean” region of the rod has little-to-no carbon coating, whereas the “deep” region has the thickest coating of carbon (Goebel et al., 1999; Goebel, 2000)). In our PIC simulations, using CST Studio Suite, we scale the loss tangent of the dielectric rods between $\tan\left(\delta\right)$ and $2000\tan\left(\delta\right)$, corresponding to $\alpha_{\mathrm{min}}$ and $\alpha_{\mathrm{max}}$, respectively, using the position-dependent Gaussian profile mentioned above. The loss tangent values used in CST Studio Suite are sampled at a single frequency of 9.3 GHz, see Sec. IV. In the CST simulations, the loss tangent at each position is assumed to be constant with respect to frequency, which is a reasonable approximation over the desired operating frequency range.
Figure 5: Slow-wave structure, emission source, electron beam, collector, and pin-to-helix transitions for a single-stage helix TWT. Electron bunches become denser along the SWS length due to beam-wave interaction.
@@
For our example, the scattering parameters of the cold single-stage TWT, determined from the time-domain solver of CST Studio Suite, indicate a reflection coefficient below -10 dB and a transmission coefficient well below -20 dB (due to attenuation) over the chosen operating frequency range, as shown in Fig. 6, allowing the input and output ports to be well-matched with good isolation to mitigate backward wave and/or regenerative oscillations. The discrepancies between the $S_{21}$ of our TL model (with the coupling coefficient $a=0$) and the simulated $S_{21}$(with loss patterning) in Fig. 6 are due to the fact that the TL model does not consider the effect of pin-to-helix transitions that are present in full-wave simulations.
@@ VI.2 Two-Stage TWT Transfer Matrix
Figure 7: Slow-wave structure with sever gap, emission source, electron beam, collector and pin-to-helix transitions for a two-stage helix TWT. Electron bunching occurs along the SWS in both stages and bunches are unaffected by the sever gap..
@@
Each TWT stage is discretized into $S_{1}=S_{2}=200$ homogeneous segments of lengths $\Delta l_{1}=l_{1}/S_{1}$ and $\Delta l_{2}=l_{2}/S_{2}$, respectively. The scattering parameters of the two-stage cold TWT are shown in Fig. 8, which indicate an adequate match at the input and output ports, as well as adequate isolation between the ports. Again, the discrepancies between the scattering parameters of our TL model (with the coupling coefficient $a=0$) and the full-wave simulations of the scattering parameters (with loss patterning) in Fig. 8 are due to the fact that the TL model does not consider the effect of pin-to-helix transitions that are present in full-wave simulations. We obtain an equivalent system transfer matrix for the input and output stages of the TWT, denoted by subscripts ST1 and ST2 respectively, by cascading the homogeneous transfer matrices $\underline{\boldsymbol{\mathrm{T}}}_{1,s}$ and $\underline{\boldsymbol{\mathrm{T}}}_{2,s}$ computed from Eqn. (4) in the forward direction (left multiplied) for each discrete position in each TWT stage for the input stage (bunching stage) as
@@
However, even though an attenuation pattern is typically used to provide isolation between TWT stages, helix severs are still non-ideal; they reflect a large portion of the RF wave (which is typically attenuated by loss patterning) and allow a small portion of the guided wave to traverse the sever gap due to capacitive coupling between helices. Thus, in our model, the RF transmission lines of stage 1 and 2 are connected using a two-port capacitive pi network with an ABCD matrix of $\underline{\boldsymbol{\mathrm{T}}}_{\mathrm{gap},\mathrm{c}}$, which describes the propagation of guided EM waves through the sever gap. The formulation of $\underline{\boldsymbol{\mathrm{T}}}_{\mathrm{gap},\mathrm{c}}$ is provided in detail in Appendix D. Two port capacitive networks have also been used to represent center conductor discontinuities in coaxial transmission lines (Whinnery, Jamieson, and Robbins, 1944; Green, 1965; Dawirs, 1969), which is a similar geometric configuration to the sever problem studied here.
@@
that models how the space charge waves along the electron beam behave in the drift-space of the sever gap. In the sever gap, there should be no coupling between the space charge wave and the EM guided mode, since there is no helix in the gap. From the beam equations in (Rouhi et al., 2021), we define the sever gap system matrix for the beam as
@@ VIII.1 Single-Stage TWT
The single-stage TWT example was simulated using the PIC solver of CST Studio Suite with a simulation duration of 10 ns, approximately 21.3 million mesh cells, and approximately 1.86 million macroparticles. Also, we use PIC simulations to confirm that the TWT is zero-drive stable. In this example, the SWS of the TWT consists of $N_{\mathrm{c}}=95$ unit cells. The peak gain of our model (with perfect matching at the input and output ports) is within approximately 0.9 dB of the peak gain found using PIC simulations. However, the frequency of the peak gain is shifted by approximately 0.4 GHz, as shown in Fig. 11a. Interestingly, the gain from LATTE and our model agree well below approximately 10 GHz, but the frequency and amplitude of the peak gain of our model agree better with CST PIC simulations, especially at frequencies above the frequency of peak gain (above 12 GHz). Using our model, the state vector was sampled at discrete positions along the TWT after being cascaded through transfer matrices, as illustrated in Fig. 3. The power of the RF wave guided along the circuit, $P(z)=\frac{1}{2}\mathrm{Re}\left(V(z)I^{*}(z)\right)$ , is plotted vs distance at an operating frequency of 12 GHz, with and without a $-10\,\mathrm{dB}$ mismatch at the TWT ports, as shown in Fig. 11b. Plotting the RF power against the position-dependent attenuation profile of the SWS, we observe that the RF power increases for a short distance into the Gaussian attenuation profile, drops close to the initial input power level near the attenuation peak, and then grows approximately exponentially with distance near the output end of the TWT. Note that the small dip in RF power on the SWS from 0 mm to 15 mm can be attributed to the “launching losses” of Pierce theory (Pierce, 1951), where the guided EM wave deposits power to the space charge wave to modulate it at the start of the tube.
@@
When mismatches are present at the input and output ports of the TWT, it appears that the power of the RF wave at the output end of the TWT is larger than the RF output power for the matched TWT in Fig. 11b, but the gain plots with and without mismatches in Fig. 11a are quite similar. This apparent discrepancy is due to the fact that our power gain definition is the transducer power gain, which depends on the available input power for the terminated TWT, rather than the constant $-10\,\mathrm{dBm}$ accepted (i.e., input) power used in our examples. When there is a mismatch at the input port of the TWT, the available RF power $P^{\mathrm{i}}=V_{\mathrm{S}}^{2}/(8Z_{\mathrm{S}})$ will be higher than the accepted RF power. The RF output power delivered to the load impedance at the end of the TWT, however, is still calculated as $P^{\mathrm{o}}=\frac{1}{2}\mathrm{Re}\left(V^{\mathrm{o}}I^{\mathrm{o}*}\right)$ and is consistent with the RF power shown in Fig. 11b.
@@
Figure 11: (a) Gain vs frequency for a single-stage TWT with no mismatches on both ports ($\Gamma=0$, dashed blue line), and with a $-10\,\mathrm{dB}$ reflection on both ports ($\Gamma=0.1$, solid blue line) calculated using model described in this paper and compared to PIC and LATTE simulations. (b) Plot of attenuation coefficient and calculated RF power along the SWS versus longitudinal position at 12 GHz with and without a $-10\,\mathrm{dB}$ mismatch (dashed red and solid red lines, respectively) at the TWT ports using our generalized Pierce model.

### v2

@@ V Interaction Impedance and Attenuation Coefficient of the SWS
where $\beta_{\mathrm{c}}$ is the phase constant within the fundamental Brillouin zone, defined here as $\beta_{\mathrm{c}}d/\pi\in\left[-1,1\right]$. For a helix-type TWT, the interaction impedance is typically evaluated within the fundamental Brillouin zone, where the beam-wave interaction is desired to occur. However, for space harmonic SWSs, such as the serpentine waveguide or coupled-cavity structure, one must consider higher order space harmonics when determining interaction impedance, like for the serpentine-type TWT studied in (Rouhi et al., 2024). Furthermore, $\left|E_{z}\left(\beta_{\mathrm{c}}\right)\right|$ is the magnitude of the axial electric field phasor (i.e., the $z$-component), for a given phase constant, along the center of the SWS where the electron beam will be introduced, and $P\left(\beta_{\mathrm{c}}\right)$ is the time-average total power flux of the mode propagating through the SWS (Gewartowski and Watson, 1965, (Ch. 10)). The quantity $\left|E_{\mathrm{z}}\left(\beta_{\mathrm{c}}\right)\right|$ is calculated from full-wave eigenmode simulations (MM, Aditya, and Chua, 2017; de Alleluia et al., 2020; Marosi et al., 2022) as
@@
Conveniently, the EM energy simulated within the unit cell between periodic boundaries in the eigenmode solver of CST Studio Suite is always normalized to 1 Joule for each eigenmode solution. Thus, the time average power flux may be simply calculated as $P=\left(1\leavevmode\nobreak\ \mathrm{Joule}\right)v_{\mathrm{g}}/d$, where the group velocity $v_{\mathrm{g}}=d\omega/d\beta_{\mathrm{c}}$ can be determined directly from the dispersion diagram using numerical differentiation. Note that the eigenmode solver of CST Studio Suite neglects losses when finding the wavenumbers and EM fields of eigenmodes, so the electric field and wavenumbers used for computing the interaction impedance also neglect the effect of losses.
@@ VI.1 Single-Stage TWT Transfer Matrix
Our first example structure, the single stage TWT, is composed of a uniform-pitch helix SWS with $N_{\mathrm{c}}=95$ turns (SWS length $l=N_{\mathrm{c}}d=98.8$ mm), with unit cell dimensions shown in Fig. 1 and described in Section II. Note again that, in our examples, the number of unit cells, $N_{\mathrm{c}}$, is different from the number of segments, $S$, that the TWT stage is divided into to make the model more flexible. Based on the specific loss profiles shown in Appendix C, we implement a nonuniform position-dependent Gaussian-shaped loss pattern in our full-wave PIC simulations by setting a loss tangent for the dielectric support rods in each individual segment (with segment length $\Delta l$, as described above) along the length of the SWS, as explained in Appendix C. At the port-ends of the SWS, the loss pattern for the single-stage SWS has a frequency-dependent attenuation coefficient of $\alpha_{\mathrm{c}}=\alpha_{\mathrm{min}}$ (calculated using Eqn. (7)) . At the midpoint between the input and output ports (i.e., at $z=l/2$) the SWS has a frequency-dependent peak attenuation coefficient $\alpha_{\mathrm{max}}=80\alpha_{\mathrm{min}}$ (determined by scaling the dielectric loss tangent by a dimensionless factor of $\tau=2000$) , and an effective loss pattern length of $l_{\alpha}=30$ mm, equal to the FWHM of the Gaussian function, as is explained in Appendix C. The quantities $\alpha_{\mathrm{min}}$ and $\alpha_{\mathrm{max}}$ correspond the attenuation coefficients in the “clean” and “deep” regions of the loss-coated dielectric support rods, respectively (i.e., the “clean” region of the rod has little-to-no carbon coating, whereas the “deep” region has the thickest coating of carbon (Goebel et al., 1999; Goebel, 2000)). In our PIC simulations, using CST Studio Suite, we scale the loss tangent of the dielectric rods between $\tan\left(\delta\right)$ and $2000\tan\left(\delta\right)$, corresponding to $\alpha_{\mathrm{min}}$ and $\alpha_{\mathrm{max}}$, respectively, using the position-dependent Gaussian profile mentioned above. The loss tangent values used in CST Studio Suite are sampled at a single frequency of 9.3 GHz, see Sec. IV. In the CST simulations, the loss tangent at each position is assumed to be constant with respect to frequency, which is a reasonable approximation over the desired operating frequency range.
Figure 5: Slow-wave structure, emission source, electron beam, collector, and pin-to-helix transitions for a single-stage helix TWT. Electron bunches become denser along the SWS length due to beam-wave interaction. Reference planes for the TL model are indicated at $z=0$ and at $z=l$, between the two pin-to-helix connection points.
@@
For our example, the scattering parameters of the cold single-stage TWT, determined from the time-domain solver of CST Studio Suite, indicate a reflection coefficient below -10 dB and a transmission coefficient well below -20 dB (due to attenuation) over the chosen operating frequency range, as shown in Fig. 6, allowing the input and output ports to be well-matched with good isolation to mitigate backward wave and/or regenerative oscillations. The discrepancies between the $S_{21}$ of our TL model (with the coupling coefficient $a=0$ in the system matrix of each segment) and the simulated $S_{21}$(with loss patterning) in Fig. 6 are due to the fact that the TL model does not consider the effect of pin-to-helix transitions that are present in full-wave simulations. The TL model describes the interaction region between reference plane 1 that exists just after the pin-to-helix connection at the input port and reference plane 2 just before the pin-to-helix-connection at the output port, as illustrated in Fig. 5.
@@ VI.2 Two-Stage TWT Transfer Matrix
Figure 7: Slow-wave structure with sever gap, emission source, electron beam, collector and pin-to-helix transitions for a two-stage helix TWT. Electron bunching occurs along the SWS in both stages and bunches are unaffected by the sever gap. Reference planes for the TL model are indicated at $z=0$ and at $z=l_{1}+l_{\mathrm{gap}}+l_{2}$, between the two pin-to-helix connection points.
@@
Each TWT stage is discretized into $S_{1}=S_{2}=200$ homogeneous segments of lengths $\Delta l_{1}=l_{1}/S_{1}$ and $\Delta l_{2}=l_{2}/S_{2}$, respectively. The scattering parameters of the two-stage cold TWT are shown in Fig. 8, which indicate an adequate match at the input and output ports, as well as adequate isolation between the ports. Again, the discrepancies between the scattering parameters of our TL model (with the coupling coefficient $a=0$) and the full-wave simulations of the scattering parameters (with loss patterning) in Fig. 8 are due to the fact that the TL model does not consider the effect of pin-to-helix transitions that are present in full-wave simulations, as was also explained in the single-stage example. The input and output reference planes for the TL model are illustrated in Fig. 7 and are adjacent to the pin-to-helix connection points. We obtain an equivalent system transfer matrix for the input and output stages of the TWT, denoted by subscripts ST1 and ST2 respectively, by cascading the homogeneous transfer matrices $\underline{\boldsymbol{\mathrm{T}}}_{1,s}$ and $\underline{\boldsymbol{\mathrm{T}}}_{2,s}$ computed from Eqn. (4) in the forward direction (left multiplied) for each discrete position in each TWT stage for the input stage (bunching stage) as
@@
However, even though an attenuation pattern is typically used to provide isolation between TWT stages, helix severs are still non-ideal; they reflect a large portion of the RF wave (which is typically attenuated by loss patterning) and allow a small portion of the guided wave to traverse the sever gap due to capacitive coupling between helices. Thus, in our model, the RF transmission lines of stage 1 and 2 are connected using a two-port capacitive pi network with an ABCD matrix of $\underline{\boldsymbol{\mathrm{T}}}_{\mathrm{gap},\mathrm{c}}$, which describes the propagation of guided EM waves through the sever gap. The formulation of $\underline{\boldsymbol{\mathrm{T}}}_{\mathrm{gap},\mathrm{c}}$ is provided in detail in Appendix D. Two port capacitive networks have also been used to represent center conductor discontinuities in coaxial transmission lines (Whinnery, Jamieson, and Robbins, 1944; Green, 1965; Dawirs, 1969), which is a similar geometric configuration to the sever problem studied here. Note that, in the sever gap region, the RF wave is expected to be quite weak due to heavy distributed attenuation on either side of the sever gap. The resistive losses that can exist in the sever gap region are only due to RF wall currents and are neglected here because they are usualy significantly smaller than the total losses on either side of the sever gap, and only capacitive coupling betwen helix stages is considered.
@@
that models how the space charge waves along the electron beam behave in the drift-space of the sever gap, like what was done in (Wong et al., 2018). In the sever gap, there should be no coupling between the space charge wave and the EM guided mode, since there is no helix in the gap. From the beam equations in (Rouhi et al., 2021), we define the sever gap system matrix for the beam as
@@ VIII Results: Gain and Evolution of the State Vector
The input power of $-10\leavevmode\nobreak\ \mathrm{dBm}$ was selected to ensure linear operation for the TWT in PIC simulations, since our model builds on linearized small-signal equations like in conventional Pierce theory (Pierce, 1947, 1951, 1950). We have verified in PIC simulations that an input power of $-10\leavevmode\nobreak\ \mathrm{dBm}$ drives the TWT well below the 1 dB gain compression point (Guillermo, 1984, Ch. 4) for both TWT examples in this study. Larger input powers on the order of $30\leavevmode\nobreak\ \mathrm{dBm}$ or higher, for example, may drive the TWT into the nonlinear saturation regime in reality and in full-wave PIC simulations; in the saturation regime, our small-signal model cannot accurately reproduce the results of full-wave PIC simulations and large-signal TWT codes such as LATTE should be used.
@@ VIII.1 Single-Stage TWT
The single-stage TWT example was simulated using the PIC solver of CST Studio Suite with a simulation duration of 10 ns, approximately 21.3 million mesh cells, and approximately 1.86 million macroparticles. Also, we use PIC simulations to confirm that the TWT is zero-drive stable. In this example, the SWS of the TWT consists of $N_{\mathrm{c}}=95$ unit cells. The peak gain of our model (with perfect matching at the input and output ports) is within approximately 0.9 dB of the peak gain found using PIC simulations. However, the frequency of the peak gain is shifted by approximately 0.4 GHz, as shown in Fig. 11a. Interestingly, the gain from LATTE and our model agree well below approximately 10 GHz, but the frequency and amplitude of the peak gain of our model agree better with CST PIC simulations, especially at frequencies above the frequency of peak gain (above 12 GHz). Using our model, the state vector was sampled at discrete positions along the TWT after being cascaded through transfer matrices, as illustrated in Fig. 3.
The power of the RF wave guided along the circuit, $P(z)=\frac{1}{2}\mathrm{Re}\left(V(z)I^{*}(z)\right)$ , is plotted vs distance at an operating frequency of 12 GHz, with and without a $-10\,\mathrm{dB}$ mismatch at the TWT ports with our analytic model, as shown in Fig. 11b. Additionally, we plot the approximate RF power of the guided wave along the TWT using data from electric field monitors in PIC simulations. These monitors are placed periodically along the center of the beam in the TWT (with period $d$), between each unit cell. The approximate RF power using these field monitors is calculated from Eqn. (5) using the steady-state ac magnitude of the longitudinal electric field (rather than the fundamental spatial harmonic of $E_{z}$) as $P_{\mathrm{PIC}}\approx\left|E_{z}\right|^{2}/\left(2\beta_{c}^{2}Z_{\mathrm{P}}\right)$. The approximate RF power found from PIC field monitors is reasonably close to the RF power calculated by our model at the collector-end of the TWT for two reasons: (i) the fundamental space harmonic of $E_{z}$ carries more energy than any other space harmonic and is close to in magnitude to the average $E_{z}$ integrated over the unit cell period in eigenmode simulations for the helix SWS, (ii) the growing wave becomes dominant over the other three wave solutions near the collector-end of the TWT (Pierce, 1951, 1950). Note that the small dip in RF power on the SWS from 0 mm to 15 mm can be attributed to the “launching losses” of Pierce theory (Pierce, 1951, 1950), where the guided EM wave deposits power to the space charge wave to modulate it at the start of the tube.
@@
When mismatches are present at the input and output ports of the TWT, it appears that the power of the RF wave at the output end of the TWT is larger than the RF output power for the matched TWT in Fig. 11b, but the gain plots with and without mismatches in Fig. 11a are quite similar. This apparent discrepancy is due to the fact that our power gain definition is the transducer power gain, which depends on the available input power for the terminated TWT, rather than the constant $-10\,\mathrm{dBm}$ accepted (i.e., input) power used in our examples. When there is a mismatch at the input port of the TWT, the available RF power $P^{\mathrm{i}}=V_{\mathrm{S}}^{2}/(8Z_{\mathrm{S}})$ will be higher than the accepted RF power. The RF output power delivered to the load impedance at the end of the TWT, however, is still calculated as $P^{\mathrm{o}}=\frac{1}{2}\mathrm{Re}\left(V^{\mathrm{o}}I^{\mathrm{o}*}\right)$ and is consistent with the RF power shown in Fig. 11b. Furthermore, in Appendix G, we show a modified example of a spatially homogeneous single-stage TWT with only intrinsic losses (i.e. losses associated to $\alpha_{\mathrm{min}}$) to demonstrate that the RF power vs position and gain vs frequency plots have a similar but not identical profiles to those found through the conventional three-wave theory used by Pierce for spatially uniform TWTs (which also neglects reflections and backward wave) (Pierce, 1950, Ch. 9)(Tsimring, 2006, Ch. 8).
@@
Figure 11: (a) Gain vs frequency for a single-stage TWT with no mismatches on both ports ($\Gamma=0$, dashed blue line), and with a $-10\,\mathrm{dB}$ reflection on both ports ($\Gamma=0.1$, solid blue line) calculated using model described in this paper and compared to PIC and LATTE simulations. (b) Plot of attenuation coefficient and calculated RF power along the SWS versus longitudinal position at 12 GHz with and without a $-10\,\mathrm{dB}$ mismatch (dashed red and solid red lines, respectively) at the TWT ports using our generalized Pierce model, in addition to the approximate RF power calculated from field monitors in PIC simulations (black line with dots).